OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (-5, 9). y = -5 y = 9 x = -5 x = 9

OpenStudy (anonymous):

Horizontal lines, are straight going left to right, aka the slope is 0. There are two ways to finish this problem, make the equation y-y(1)=m(x-x(1)) where you would get y-9=0(x-(-5)). then it would just go to y-9=0 or y=9. or just look at the point, it crosses at y=9, and there is no slope, so the y is always 9. hope that kind of explains it

OpenStudy (anonymous):

B? @sschroeder